OpenStudy (anonymous):

the average of integers a and b is 12. if c=3 what is the average of integers a b and c.

OpenStudy (anonymous):

9

OpenStudy (radar):

(a+b)/2=12 a could be 6 b could 6, a could be 9 b could be 3 etc. now we bring in c (a + b + c)/3= ??? regardles we know tha a+b=12, so a+b+c=15 15/3 = 5 The new average is 5